PLEASE SOLVE AND SHOW IN EXCEL:

A teacher would like to know if her lecture made an impact on students' learning in her class. She gives the students a pre-test before the lecture. Then after she teaches, she gives a post-test. The post-test has similar questions as the pre-test. The results are:

Pre-test

Post-test

78

80

67

69

56

70

78

79

96

96

82

84

84

88

90

92

87

92

The average of the post-test is (83.33) and the pretest average is (79.78). She wants to know if the lesson had a significant impact on student learning based on these scores. Did the students score significantly higher on the post-test than they did on the pre-test?

State your hypothesis and your null hypothesis:

Decide which t-test you will use (Dependent, Independent – equal variance, Independent – unequal variance):

Test your null hypothesis at the .05 level (decide if this is one-tailed or two-tailed). Which tailed test did you use and what is your observed t value?

What is the critical value for this test?

Do you accept or reject your null hypothesis?

State your conclusion:
Answer #1

Ans:

d=Post test-Pre test

Ho: Md < 0

H, :μα > 0

pre- test Post-test d
1 78 80 2
2 67 69 2
3 56 70 14
4 78 79 1
5 96 96 0
6 82 84 2
7 84 88 4
8 90 92 2
9 87 92 5
d-bar 3.556
sd 4.187

Dependent samples t test

One tailed test

Test statistic:

t=(3.556-0)/(4.187/sqrt(9))

t=2.548

df=9-1=8

p-value=tdist(2.548,8,1)=0.0171

critical t value=1.860

Reject the null hypothesis.

There is sufficient evidence to conclude that the students score significantly higher on the post-test than they did on the pre-test